OpenStudy (anonymous):

and i also need help to solve this one the last one by graphing just give me please the two points 3x+4y=20 3x-2y=8

OpenStudy (anonymous):

any points... it always nice to pick x=0 - find y=... and let say x=1 - find corresponding y. do the same for the second equation

OpenStudy (anonymous):

1) x=0, so y=5 and x=2, y=7/2 2) x=0, y=-4 y=0, x=8/3 you can pick ANY values of x or y & find corresponding values of y & x